OpenStudy (stefrheart):

Write an equation for the line with x-intercept 5 and y-intercept 3.

OpenStudy (deepolisnoob):

The y-intercept is going to be equal to 3, so you can assume that in your y=mx+b equation, your b is equal to 3, so now you have y=mx+3 Now you need to find a slop that will match your line. When your x-intercept is equal to 5, that means that you have a coordinate point at (5,0). Since your y-intercept is at 3, you have another point at (0,3). From here, you can find slope by calculating rise/run. To go from (0,3) to (5,0), you rise -3, and you run 5, making your slope -3/5. Your entire equation would be y=-3/5x+3

OpenStudy (stefrheart):

Thanks so much!!!!